Find the ratio of the speed of light in fused quartz to the speed of light in flint glass.
Answer #1

here,


refractive index of flint glass is n1 =(1.57- 1.75)

average value of n1=1.66 and speed of the light is v1

refractive index of fused quartz glass is n2=1.46 and speed of the light is v2


use,

refractive index n=(speed of the light in vacuum)/(speed of the light in material)

n=C/V

===>

v2/v1= n1/n2

v2/v1=1.66/1.46

v2/v1=1.14